2 of 12 (1 complete) HW Score: 5.56%, 0.67 of 12 pts Score: 0 of 1 pt P 14-2 (book/static) Question Help Starware Software waPlease answer parts A and B
Answer #1

a-1). New Total Number of shares = Founder's Shares / [1 - Ownership% desired by venture capitalist]

= 8 million / [1 - 0.20] = 8 million / 0.80 = 10 million shares

Shares Owned by Venture Capitalist = New total number of shares - Founder's shares

= 10 million - 8 million = 2 million

a-2). Implied price per share = Invested Amount by venture capitalist / Shares Owned by Venture Capitalist

= $1 million / 2 million = $0.50 per share

b). Post-money Valuation = Implied price per share * New Total Number of shares

= $0.50 * 10 million = $5 million